Sol Summary Report – March 1th

Crew 275 Sol Summary Report 01Mar2023

Sol: 17

Summary Title: Saving Private LOAC

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

The weather surprised us again today and we had to cancel our exploration EVA to Candor Chasma, initially planned in the morning. When it stopped snowing, Mission Support agreed to let us put on our spacesuits to quickly change the batteries of our atmospheric instruments and retrieve the LOAC device, as it is very sensitive to precipitation. With the rainy afternoon that we had, we were indeed happy that the LOAC was safe within the Hab!

All these changes in the schedule give additional work to our Crew Scientist, who is responsible for planning our scientific activities. At least, we all hope that the ground will soon be dry so that we can go back on EVA, especially during the visit of the BBC on Friday (Sol 19).

Today, we have finished equipping everyone with the tracking system, so that we are aware of our position at anytime of the day and correlate this with other information. We have also continued the human factors protocols, as well as the tests of the ultrasound device based on virtual reality and artificial intelligence.

Look Ahead Plan:

We would like to reschedule today’s EVA at Candor Chasma. The plan would be to have a first EVA on Friday (Sol 19) and another one on Saturday (Sol 20). The EVA crews will have to find specific checkpoints based on a 2D map or a 3D render, respectively. We would like to take the opportunity of these EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather.

The BBC is planned to arrive at the MDRS on Friday (Sol 19) between 8:30 and 9:00. They would follow us on our EVA to Candor Chasma, then film inside the station in the afternoon, and leave at around 17:00.

Anomalies in work: None.

Weather: Rainy, low 0°C/32°F, high 8°C/46°F.

Crew Physical Status: Optimal.

EVA: The objectives of EVA #14 were changed due to the bad weather conditions. This EVA will have to be rescheduled.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site